# A Model-Agnostic Graph Neural Network for Integrating Local and Global Information

### v3

@@ Abstract
Graph Neural Networks (GNNs) have achieved promising performance in a variety of graph-focused tasks. Despite their success, however, existing GNNs suffer from two significant limitations: a lack of interpretability in results due to their black-box nature, and an inability to learn representations of varying orders. To tackle these issues, we propose a novel Model-agnostic Graph Neural Network (MaGNet) framework, which is able to effectively integrate information of various orders, extract knowledge from high-order neighbors, and provide meaningful and interpretable results by identifying influential compact graph structures. In particular, MaGNet consists of two components: an estimation model for the latent representation of complex relationships under graph topology, and an interpretation model that identifies influential nodes, edges, and node features. Theoretically, we establish the generalization error bound for MaGNet via empirical Rademacher complexity, and demonstrate its power to represent layer-wise neighborhood mixing. We conduct comprehensive numerical studies using simulated data to demonstrate the superior performance of MaGNet in comparison to several state-of-the-art alternatives. Furthermore, we apply MaGNet to a real-world case study aimed at extracting task-critical information from brain activity data, thereby highlighting its effectiveness in advancing scientific research.
@@ 1 Introduction
Graph-structured data is ubiquitous throughout the natural and social sciences, from brain networks to social relationships. In the most general view, a graph is simply a collection of nodes representing entities such as people, genes, and brain regions, along with a set of edges representing interactions between pairs of nodes. By representing such interconnected entities as graphs, it is possible to leverage their geometric topology to study functional relationships in systems with network-based frameworks. For example, when studying brain function, graphs can be used to model activity relationships among neurons or brain regions. For such applications, it is essential to build relational inductive biases into machine learning methods in order to develop systems that can learn, reason, and generalize. To achieve this goal, in recent years there has been a surge in research on graph representation learning, including techniques for deep graph embedding powered by deep learning architecture, e.g., neural message passing approaches inspired by belief propagation. These advances in graph representation learning have led to new developments in various domains, including the analysis of information processing in the brain, chemical synthesis, recommender systems, and modeling social networks.
@@
Among the methods for deep graph representation, the family of graph neural networks (GNNs) has achieved remarkable success in real-world graph-focused tasks (Defferrard et al., 2016; Zhang and Chen, 2018; Ying et al., 2019; Veličković, 2023). In general, GNNs iteratively aggregate and combine node representations within a graph, through a process called message passing, to generate a set of learned hidden representation features (Gilmer et al., 2017). The main neural architectures of GNNs include graph convolutional networks (GCNs; Kipf and Welling (2016)), graph attention networks (GATs; Veličković et al. (2017)), graph transformer networks (GTNs; Yun et al. (2019)) and many other variants (Hamilton, 2020).
@@
While GNNs are capable of capturing subgraph information through local operations, e.g., graph convolution operations, they can be prone to over-smoothing the learned representations when applying multiple rounds of local operators. This can lead to models that treat all nodes uniformly, resulting in node representations that converge towards an indistinguishable vector (Li et al., 2018). Additionally, over-smoothing could limit the ability to capture high-order information, which can only be aggregated through executing a sufficient number of local operations, especially when nodes have a large receptive field (Hamilton, 2020). Several studies have suggested that this over-smoothing phenomenon is a key factor in the performance degradation of deep GNNs (Bodnar et al., 2022). To avoid this issue, some recent works (Zhang et al., 2022; Keriven, 2022) have advocated the use of various shallow GNNs (e.g., up to three layers). This, as discussed above, results in failure to capture high-order information due to insufficient rounds of message passing. In addition, Zhao and Akoglu (2019) and Yang et al. (2020) develop normalization layers to avoid the embeddings of different nodes being indistinguishable. However, the use of these normalization layers greatly increases the difficulties in model training and constrains the expressive power of GNNs.
One main premise of this paper is that to improve the overall representation power of GNNs, we need to build new learning mechanisms that directly incorporate and effectively combine information from neighbors at different orders. Statistically, by incorporating both low-order (immediate neighbors) and high-order (i.e., neighbors beyond the immediate vicinity) information, GNNs can learn a richer representation under graph topology. For example, in the brain, low-order information may capture relationships within a functional region, while high-order information captures relationships across functional regions (Bassett and Sporns, 2017). Importantly, the information flows in a sequential manner starting from within-region functionality to cross-region functionality (Vázquez-Rodríguez et al., 2019).
@@
Models that adhere to the principle of effectively combining different-order information are known as multi-scale GNNs, as they enable the exploration and integration of information at multiple levels of granularity within a graph (Gasteiger et al., 2018; Xu et al., 2018b; Liao et al., 2019; Sun et al., 2019; Oono and Suzuki, 2020; Liu et al., 2022). Multi-scale GNNs provide enhanced flexibility and adaptability in facilitating information flow across various neighborhood ranges. This is achieved by directing the outputs of intermediate layers to contribute to the final representation. Existing methods however struggle to effectively integrate representations of different orders in a sequential manner due to their memoryless property: at each GNN layer, node representations are updated entirely based on the current input from their immediate neighbors, without directly retaining information from previous layers. In other words, the most recently learned representation tends to modify the previously acquired representation and cause information loss from previous layers. Furthermore, this approach still suffers from over-smoothing issues and generally fails to capture high-order latent representations.
To address these issues, we develop a novel Model-agnostic Graph neural Network (MaGNet) framework with two components: the estimation model and the interpretation model. The estimation model captures the complex relationship between the feature information and a target outcome under graph topology, allowing for powerful latent representation. The interpretation model, on the other hand, identifies a compact subgraph structure, i.e., influential nodes and edges, along with a small subset of node features that have a crucial role in the learned estimation model. This provides a meaningful interpretation of the underlying mechanisms of the relationships, allowing for a better understanding of the factors governing the behavior of complex dynamic systems. As an example, for understanding brain dynamics, the interpretation model can decode the most informative neural dynamics.
The main advantages of the proposed framework and our contributions are summarized as follows. First, the proposed neural architecture of the model alleviates the over-smoothing issue and thus can effectively extract knowledge from high-order neighbors. In addition, the proposed actor-critic neural architecture has the ability to effectively integrate information of various orders via resolving the memoryless issue. To this end, it adaptively combines the representation learned from a set of actor graph neural networks, each attending to a specific information order, while a critic neural network evaluates the quality of the hidden representation learned by each actor network.
Second, we construct a practically useful interpretation framework for the black-box estimation model. We formulate our interpretation model as an optimization task that maximizes the information gained between a prediction of the estimation model and the distribution of possible subgraph structures. This approach is model-agnostic and can interpret the estimation model on any graph-focused task without assumptions regarding the underlying true model or data-generating mechanisms.
Third, we study the ability to integrate various-order information as well as the statistical complexity of the proposed model via a minimax measurement on empirical Rademacher complexity. Unlike existing analyses solely working on standard message passing neural networks, our results can be applied to a mixture of neural architectures between message passing neural networks and feedforward neural networks, where the message passing neural network is a special case in our theoretical framework. Furthermore, we provide a rigorous generalization error bound for the MaGNet estimation model that consists of sequential deep-learning component models.
The rest of the paper is organized as follows. Section 2 lays out the basic graph notation and model settings. Section 3 formally defines the estimation module. Section 4 illustrates how to construct the interpretation model for an estimated model. A comprehensive theoretical study of the proposed framework is provided in Section 5. Sections 6 and 7 demonstrate the empirical performance of our methods. We conclude this paper with a discussion of possible future research directions. All technical proofs are provided in the Appendix.
@@ 3.1 Actor-Critic Graph Neural Network
Further, in contrast to node classification, our interest lies in graph classification tasks. Therefore, based on the learned node representation $H^{(k)}$, we can apply a graph pooling operation to summarize the graph embedding from $H^{(k)}$. The goal of the graph pooling operation is to aggregate information across the entire graph to produce a single, fixed-size representation. Specifically, we could use a simple average graph pooling operation to obtain graph embeddings with a compact node representation: $\bar{H}^{(k)}=\frac{1}{N}\mathbf{1}_{N}H^{(k)}$ See the Appendix for other types of graph pooling operation, and Hamilton (2020) for a more comprehensive review.
@@ 4 Interpretation Model
Essentially, information gain quantifies the change in prediction probability between the full model $p_{\widehat{\theta}}(\cdot)$ and the one constrained to the subgraph $G_{sub}$ and the subset node feature $X_{sub}$. For example, if removing edge $e_{ij}$, i.e., the $(i,j)$th element in the adjacency matrix $A$, from the full graph $G$ significantly decreases the prediction probability, then this edge is influential and should be included in the subgraph $G_{sub}$. Conversely, if the edge $e_{ij}$ is deemed redundant for prediction by the learned estimation model, it should be excluded.
@@
where $G_{sub}(\mu)$ is the parametrized random subgraph. Due to the discrete nature of the subgraph $G_{sub}(\mu)$, the objective function is non-smooth, making optimization challenging and unstable. To address this issue, we further propose a continuous approximation for the binary sampling process (Maddison et al., 2016). Let $\epsilon$ be a uniform random variable, i.e., $\epsilon\sim\text{Unif}(0,1)$, and the real-valued parameters $\psi_{ij}\in\Psi$, and a temperature parameter $\omega\in\mathbb{R}^{+}$, then a sample of the binary edge $e_{ij}$ can be approximated by a sigmoid mapping:
@@
We denote $\widetilde{G}_{sub}(\Psi)$ as the continuous relaxation counterpart of the subgraph, with the $(i,j)$th element of the adjacency matrix being $\widetilde{e}_{ij}$. Interestingly, in our analysis, the temperature parameter $\omega$ can describe the relationship between $\widetilde{G}_{sub}(\Psi)$ and $G_{sub}(\mu)$. We observe that as $\omega\rightarrow 0$, the approximated edge $\widetilde{e}_{ij}$ converges to the edge $e_{ij}$, with the probability mass function, $\lim_{\omega\rightarrow 0}P(\widetilde{e}_{ij}=1)=\frac{\exp(\psi_{ij})}{1+\exp(\psi_{ij})}.$ Recall that the edge $e_{ij}$ follows a Bernoulli distribution with mean $\mu_{ij}$. If we reparameterize $\psi_{ij}$ such that $\psi_{ij}=\log\left(\frac{\mu_{ij}}{1-\mu_{ij}}\right)$, it achieves asymptotical consistency of the approximated subgraph by following the limiting theory (Paulus et al., 2020), i.e., $\lim_{\omega\rightarrow 0}\widetilde{G}_{sub}(\Psi)={G}_{sub}(\mu)$. This supports the feasibility of applying continuous relaxation to the binary distribution.
@@ 5 Theory
The above assumptions are common in the (graph) neural network literature. Assumptions 5.1-5.2 impose norm constraints on the parameters and input feature, making the model class fall into a compact metric space (Liao et al., 2020). Assumption 5.3 is a standard assumption to control the intensity of the graph Laplacian in GNN literature (Hamilton, 2020). Assumption 5.4 requires us to focus on homogeneous graphs (Liao et al., 2020; Lv, 2021). Assumption 5.5 is a standard assumption in bounding the width of neural network layers.

### v4

@@ Abstract
Graph Neural Networks (GNNs) have achieved promising performance in a variety of graph-focused tasks. Despite their success, however, existing GNNs suffer from two significant limitations: a lack of interpretability in their results due to their black-box nature, and an inability to learn representations of varying orders. To tackle these issues, we propose a novel Model-agnostic Graph Neural Network (MaGNet) framework, which is able to effectively integrate information of various orders, extract knowledge from high-order neighbors, and provide meaningful and interpretable results by identifying influential compact graph structures. In particular, MaGNet consists of two components: an estimation model for the latent representation of complex relationships under graph topology, and an interpretation model that identifies influential nodes, edges, and node features. Theoretically, we establish the generalization error bound for MaGNet via empirical Rademacher complexity, and demonstrate its power to represent layer-wise neighborhood mixing. We conduct comprehensive numerical studies using simulated data to demonstrate the superior performance of MaGNet in comparison to several state-of-the-art alternatives. Furthermore, we apply MaGNet to a real-world case study aimed at extracting task-critical information from brain activity data, thereby highlighting its effectiveness in advancing scientific research.
@@ 1 Introduction
Graph-structured data is ubiquitous throughout the natural and social sciences, from brain networks to social relationships. A graph is simply a collection of nodes representing entities such as people, genes, and brain regions, along with a set of edges representing interactions between pairs of nodes. By representing such interconnected entities as graphs, it is feasible to leverage their geometric topology to study statistical relationships among nodes using network-based frameworks. Among graph representation methods, the family of graph neural networks (GNNs) has achieved remarkable success in real-world graph-based tasks (Veličković, 2023). In general, GNNs iteratively aggregate and combine node representations within a graph, through a process called message passing, to generate a set of learned hidden representation features. The main neural architectures of GNNs include graph convolutional networks (GCNs; Kipf and Welling, 2016), graph attention networks (GATs; Veličković et al., 2017), graph transformer networks (GTNs; Yun et al., 2019), among many other variants.
@@
While GNNs are capable of capturing subgraph information through message passing, they can be prone to over-smoothing the learned representations when applying multiple rounds of message passing operations. This can cause models to treat all nodes uniformly, leading to node representations that converge into indistinguishable entities (Li et al., 2018). Additionally, over-smoothing could limit the ability to capture high-order information, which can be only aggregated through a sufficient number of message passing operations (Hamilton, 2020). Several studies suggest that over-smoothing significantly contributes to deep GNN performance degradation (Bodnar et al., 2022). To address this issue, (Zhang et al., 2022) advocated for using shallow GNNs (e.g., up to three layers). However, this approach fails to capture high-order information due to insufficient message passing. Additionally, Zhao and Akoglu (2019) and Yang et al. (2020) developed normalization layers to prevent node embeddings from becoming indistinguishable, but this increases training difficulty and limits the expressive power of GNNs.
@@
A main premise of this paper is that to enhance the overall representation power of GNNs with statistical guarantees, we need to develop new learning mechanisms that directly incorporate and effectively combine information from neighbors at different orders. Statistically, by integrating both low-order (i.e., immediate neighbors) and high-order (i.e., neighbors beyond the immediate vicinity) information, GNNs can learn a richer and more complete representation under the graph topology. Models that follow the principle of effectively combining information from different orders are known as multi-scale GNNs, as they enable the exploration and integration of information at different levels of granularity within a graph (Xu et al., 2018b; Sun et al., 2019; Oono and Suzuki, 2020; Liu et al., 2022). The main idea is to direct the outputs of intermediate layers to contribute to the final representation. Existing methods however struggle to effectively integrate representations of different orders in a sequential manner due to their memoryless property: at each GNN layer, node representations are updated entirely based on the current input from their immediate neighbors, without directly retaining information from previous layers. Furthermore, Multi-scale GNNs still suffer from over-smoothing issues and generally fail to capture high-order latent representations.
To address the aforementioned issues, we develop a novel Model-agnostic Graph neural Network (MaGNet) framework consisting of two components: the estimation model and the interpretation model. The estimation model captures the complex relationship between the feature information and a target outcome under graph topology, allowing for powerful latent representation corresponding to unique-order graph information. The interpretation model, on the other hand, identifies a compact subgraph structure – specifically, influential nodes and edges – along with a small subset of node features that play a crucial role in the learned estimation model. The main advantages of the proposed framework, along with our contributions, are outlined below. First, the proposed neural architecture of the model alleviates the over-smoothing issue and thus can effectively extract knowledge from high-order neighbors. The proposed actor-critic neural architecture effectively integrates multi-order information by resolving the memoryless issue. It adaptively combines representations from actor graph neural networks, each focused on a specific order, while a critic network evaluates the quality of the learned representations. Second, we develop an interpretation framework, formulated as an optimization task that maximizes information gain over the distribution of possible subgraph structures. This approach is model-agnostic as there is no assumption on the true statistical models or data-generating mechanisms. Third, we study the ability to integrate various-order information as well as the statistical complexity of the proposed model via an empirical Rademacher complexity. Unlike existing analyses limited to standard message passing neural networks, our results are applicable to a mix of message passing and feedforward neural networks, with message passing networks being a special case in our framework. Furthermore, we provide a statistical generalization error bound for the MaGNet estimation model that consists of sequential deep-learning component models.
@@ 3.1 Actor-Critic Graph Neural Network
Further, in contrast to node classification, our interest lies in graph classification tasks. Therefore, based on the learned node representation $H^{(k)}$, we can apply a graph pooling operation to summarize the graph embedding from $H^{(k)}$. The goal of the graph pooling operation is to aggregate information across the entire graph to produce a single, fixed-size representation. Specifically, we could use a simple average graph pooling operation to obtain graph embeddings with a compact node representation: $\bar{H}^{(k)}=\frac{1}{N}\mathbf{1}_{N}H^{(k)}.$ See the Appendix for other types of graph pooling operation, and Hamilton (2020) for a more comprehensive review.
@@ 4 Interpretation Model
Essentially, information gain can quantify the change in prediction probability between the full model $p_{\widehat{\theta}}(\cdot)$ and the one constrained to the subgraph $G_{sub}$ and the subset node feature $X_{sub}$ Ying et al. (2019). For example, if removing edge $e_{ij}$, i.e., the $(i,j)$th element in the adjacency matrix $A$, from the full graph $G$ significantly decreases the prediction probability, then this edge is influential and should be included in the subgraph $G_{sub}$. Conversely, if the edge $e_{ij}$ is deemed redundant for prediction by the learned estimation model, it should be excluded.
@@
where $G_{sub}(\mu)$ is the parametrized random subgraph. Due to the discrete nature of the subgraph $G_{sub}(\mu)$, the objective function is non-smooth, making optimization challenging and unstable. To address this issue, we further leverage a continuous approximation for the binary sampling process (Maddison et al., 2016; Luo et al., 2020). Let $\epsilon$ be a uniform random variable, i.e., $\epsilon\sim\text{Unif}(0,1)$, and the real-valued parameters $\psi_{ij}\in\Psi$, and a temperature parameter $\omega\in\mathbb{R}^{+}$, then a sample of the binary edge $e_{ij}$ can be approximated by a sigmoid mapping:
@@
We denote $\widetilde{G}_{sub}(\Psi)$ as the continuous relaxation counterpart of the subgraph, with the $(i,j)$th element of the adjacency matrix being $\widetilde{e}_{ij}$. Interestingly, the temperature parameter $\omega$ can describe the relationship between $\widetilde{G}_{sub}(\Psi)$ and $G_{sub}(\mu)$. We observe that as $\omega\rightarrow 0$, the approximated edge $\widetilde{e}_{ij}$ converges to the edge $e_{ij}$, with the probability mass function, $\lim_{\omega\rightarrow 0}P(\widetilde{e}_{ij}=1)=\frac{\exp(\psi_{ij})}{1+\exp(\psi_{ij})}.$ Recall that the edge $e_{ij}$ follows a Bernoulli distribution with mean $\mu_{ij}$. If we reparameterize $\psi_{ij}$ such that $\psi_{ij}=\log\left(\frac{\mu_{ij}}{1-\mu_{ij}}\right)$, it achieves asymptotical consistency of the approximated subgraph by following the limiting theory (Paulus et al., 2020), i.e., $\lim_{\omega\rightarrow 0}\widetilde{G}_{sub}(\Psi)={G}_{sub}(\mu)$. This supports the feasibility of applying continuous relaxation to the binary distribution.
@@ 5 Theory
The above assumptions are common in the (graph) neural network literature. Assumptions 5.1-5.2 impose norm constraints on the parameters and input feature, making the model class fall into a compact metric space (Liao et al., 2020). In general, Assumption 5.1 does not require a specific data distribution for the feature vector. It holds as long as the feature vector has a bounded $L_{2}$-norm, regardless of its distribution. For example, feature vectors following the truncated Gaussian distribution, uniform distribution, logarithmic distribution, and autoregressive distribution within the $L_{2}$-ball all satisfy Assumption 5.1. Assumption 5.3 is a standard assumption to control the intensity of the graph Laplacian in GNN literature (Hamilton, 2020). Assumption 5.4 requires us to focus on homogeneous graphs (Liao et al., 2020; Lv, 2021). Assumption 5.5 is a standard assumption in bounding the width of neural network layers.